# Dependency equilibria: Boundary cases and their real algebraic geometry

### v1

@@ 2. Preliminaries, explanation of the results and examples
Explanation 2.10 (concerning Theorem 3.18 & Corollary 3.19).
For generic games (see Explanation 2.7), every NE is a DE by Corollary 3.19. What is more, for arbitrary games, every NE lies on the Spohn variety which is the algebraic model for the set of DE of a game (Theorem 3.18).
Figure 3 shows the Spohn variety consisting of four lines for the game
$\begin{pmatrix}-1&-1\\ -2&3\end{pmatrix}\text{ and }\begin{pmatrix}-3&0\\ 0&0\end{pmatrix}.$
The lower bound $\overline{\mathcal{V}\setminus\mathcal{W}}^{\text{Zar}}$ for DE equals the blue component. By Example 4.7, the blue point is a NE but not a dependency equilibrium. This is only possible because it lies on one of the green components of $\mathcal{V}$. These are exactly the components of $\mathcal{V}$ that are contained in $\mathcal{W}$.
Figure 3. The Spohn variety of the game from Explanation 2.10. The blue point is a NE but not a DE. The blue line goes along one of the faces of the tetrahedron. One of the three green lines passes along an edge.
Explanation 2.11 (concerning Proposition 3.21).
The proposition gives a criterion, in terms of linear algebra, for the existence of a whole family of totally mixed DE around a pure strategy. This criterion also serves as a sufficient condition for the existence of a pure DE.
The criterion is based on the Jacobian matrix of the Spohn variety at a pure strategy, which is analysed in Remark 3.20. Clearly, it is only of theoretic value without further investigations of the Jacobian. But, in principle, this can be done, as Example 4.8 shows. There, we give a very explicit form of this criterion for $(2\times 2)$-games involving only equalities and inequalities in the entries of the payoff tables. Figure 1 shows the Spohn variety of an example of a game where this criterion applies for the pure strategies $(1,0,0,0)$ and $(0,0,0,1)$, but it fails for the other two pure strategies, although these are also DE.
Explanation 2.12 (concerning Theorem 4.2).
$(2\times 2)$-games split up into distinct classes concerning the geometric structure of their sets of DE. An understanding of these classes is based on the considerations of Theorem 4.2 and its proof which give, from this viewpoint, a full and sensible characterization of $(2\times 2)$-games. Cases (3)(c) and (3)(d) deserve further investigation concerning the number of irreducible components of the Spohn variety (which all are curves in these cases) and their individual degrees. This is studied in an ongoing project [9].
3. Results for arbitrary games
Real points of the Spohn variety
In the following, we want to show that the Spohn variety $\mathcal{V}=\mathcal{V}_{X}$ of a game $X$, that is, the variety of all complex projective points satisfying the equations of a dependency equilibrium (see [10, (4)]), always has smooth points with real coordinates. This enables us to apply the following result which, in turn, makes it possible to study the set of all dependency equilibria from the point of view of complex projective varieties. As a consequence, we can apply the results of [10] to the case that is actually relevant for game-theoretical applications.
Fact 3.1.
[8, Theorem 2.2.9] Let $V$ be an irreducible complex affine variety defined by real polynomials. If $V$ contains a smooth point with real coordinates then $V(\mathbb{R})$, the set of real points of $V$, lies Zariski dense in $V$.
We can apply this by taking a dense open affine subset of the ambient projective space and thereby assuming that $V=\mathcal{V}$ is affine.
Theorem 3.2.
Proof.
For $(2\times 2)$-games, this is just Proposition 4.1.
In all other cases, we use the vector bundle parametrization $\phi:\mathcal{V}_{X}\dashrightarrow(\mathbb{P}^{1})^{n}$ of the proof of [10, Theorem 8] and can write
Regarding game-theoretic applications, it might be interesting to find smooth real points on every irreducible component of the Spohn variety of a game. The question below asks exactly for this generalization of Theorem 3.2,
Question 3.3.
The following is the combination of Theorem 3.2 with [10, Theorem 6].
Corollary 3.4.
Passing to boundary cases
The definition of DE used in [10] lets points on the relative boundary of $\Delta$ unregarded. We denote by $\overline{\Delta}$ the relative closure of $\Delta$, that is, the set of all joint strategies of the players, while $\Delta$ is just the set of totally mixed strategies.
Spohn [13, 14] defines DE as points $p=(p_{j_{1}\ldots j_{n}})$ in $\overline{\Delta}$ that are limits of sequences $(p^{(r)})_{r\in\mathbb{N}}$ in $\Delta$ satisfying
for all $i\in\{1,\ldots,n\}$ and $k,k^{\prime}\in\{1,\ldots,d_{i}\}$. In the special case of $(2\times 2)$-games with the notation of [10, Example 2], these conditions read as follows:
Spohn then claims in [14, Section 4, Observation 3] that, with this definition, every Nash equilibrium (NE) is a DE. This is true, in principle, but not in general, as we want to illustrate with the next example.
Example 3.5 (Bach or Stravinski).
We use the payoff tables $(2,0,0,1)$ and $(1,0,0,2)$ as in [14, Section 3]. The equations (1) translate as
First, we want to note that [14, Section 3] does not contain the DE of the form $(0,1-x,x,0)$ for $0\leq x\leq 1$. The two special ones with $x\in\{0,1\}$ can be reached as limits of those with $10<x<$ 1 for which the above equation hold even without limits. Second, we point out that the two NE $(1,0,0,0)$ and $(0,0,0,1)$ cannot be approximated by points in $\Delta$ satisfying the equations above. Indeed, assume to the contrary that the sequence $p^{(r)}$ converges to $(1,0,0,0)$ and the above equations are satisfied. Then, $\lim_{r\to\infty}p_{11}^{(r)}=1$ and all other $p_{ij}^{(r)}$ converge to 00. This means that $\lim_{r\to\infty}\frac{2p^{(r)}_{11}}{p^{(r)}_{11}+p^{(r)}_{12}}=2=\lim_{r\to\infty}\frac{p^{(r)}_{22}}{p^{(r)}_{21}+p^{(r)}_{22}}$. Since all the $p_{ij}^{(r)}$ are strictly positive, it follows that $\frac{1}{2}=\lim_{r\to\infty}\frac{p^{(r)}_{21}+p^{(r)}_{22}}{p^{(r)}_{22}}=\lim_{r\to\infty}\frac{p^{(r)}_{21}}{p^{(r)}_{22}}+1$. So $\lim_{r\to\infty}\frac{p^{(r)}_{21}}{p^{(r)}_{22}}=-\frac{1}{2}$ and hence this last sequence must get negative for $r$ sufficiently large. This is a contradiction to $p_{ij}^{(r)}>0$.
We now give a slightly more general definition for DE, so that every NE is a DE for generic games. Additionally, we define two further equilibrium concepts and then want to investigate the relations between all of them. The final goal would be to show (if true) that all three concepts coincide which would make DE (in our sense) computable by symbolic methods. But first, we need some notation. Let $X$ be a game given in normal form by its payoff tensors $(X^{(i)}_{j_{1}\ldots j_{n}})$, $\mathcal{V}=\mathcal{V}_{X}\subseteq\mathbb{P}^{d_{1}\times\ldots\times d_{n}-1}$ the Spohn variety of $X$ and $\mathcal{I}=\mathcal{I}_{X}$ its ideal in the polynomial ring $R=\mathbb{C}[p_{j_{1}\ldots j_{n}}\mid j_{i}\in\{1,\ldots,d_{i}\}]$. Moreover, let $\mathcal{W}\subseteq\mathbb{P}^{d_{1}\times\ldots\times d_{n}-1}$ be the union of hyperplanes defined by the polynomial
$s=\prod_{i,k}p_{+\ldots+k+\ldots+}.$
For a subset $M\subseteq\mathbb{C}^{d_{1}\times\ldots\times d_{n}}$ (respectively $\mathbb{P}^{d_{1}\times\ldots\times d_{n}-1}$ or $\mathbb{R}^{d_{1}\times\ldots\times d_{n}}$), we denote by $\overline{M}$ its usual euclidean closure, that is, the set of all points that are limits of sequences in $M$, and by $\overline{M}^{\text{Zar}}$ its Zariski closure, that is, the set of all points that are common zeros of polynomials vanishing on $M$. It is a basic fact that $\overline{M}\subseteq\overline{M}^{\text{Zar}}$ because every algebraic set is closed in the euclidean topology.
Definition 3.6.
  1. (1)
  2. (2)
    An analytic dependency equilibrium (aDE) is a point in $\overline{\Delta}\cap\overline{(\mathcal{V}\setminus\mathcal{W})}$.
  3. (3)
    A geometric dependency equilibrium (gDE) is a point in $\overline{\Delta}\cap\overline{(\mathcal{V}\setminus\mathcal{W})}^{\text{Zar}}$.
For generic games, we can replace complex tensors by real tensors in Definition 3.6 and still deduce results analogous to the contents of this section, see Remark 3.12.
@@ 3. Results for arbitrary games
  1. (1)
    Every aDE clearly is a DE because it can be approximated by a sequence in $\mathcal{V}\setminus\mathcal{W}$. In symbols, aDE $\subseteq$ DE.
  2. (2)
    Every aDE is a gDE because $\overline{(\mathcal{V}\setminus\mathcal{W})}\subseteq\overline{(\mathcal{V}\setminus\mathcal{W})}^{\text{Zar}}$. In symbols, aDE $\subseteq$ gDE.
For all $(2\times 2)$-games in [14, Section 3] (Matching pennies, Bach or Stravinski, Hawk and Dove, Prisoner’s Dilemma), the three concepts of DE, aDE and gDE coincide.
Cases where DE concide with gDE are particularly nice from the view point of game-theoretic applications of the theory of DE because we can describe DE by methods from computational commutative algebra there: Computer algebra software like Macaulay2 can compute for us the minimal prime ideals of the saturated ideal $(\mathcal{I}_{X}:s^{\infty})$. The generators of these minimal primes exactly give the equations for the irreducible components of $\overline{(\mathcal{V}\setminus\mathcal{W})}^{\text{Zar}}$. The only humanly task is then to figure out which points in $\overline{\Delta}$ satisfy these equations that are very easy in many cases. The next proposition tells us that a large class of DE always can be computed in this way.
Proposition 3.8.
For every game, aDE and gDE coincide.
Proof.
In [7, Exposé XII, Prop. 2.2], if we take $\mathcal{V}$ as the $\mathbb{C}$-scheme $X$ and $T=\mathcal{V}\setminus\mathcal{W}$ the locally constructible subset (that is actually a Zariski open subset) then this proposition just says that $\overline{\mathcal{V}\setminus\mathcal{W}}^{\text{Zar}}=\overline{\mathcal{V}\setminus\mathcal{W}}$. This immediately implies our assertion. ∎
Remark 3.9.
Theorem 3.10.
Proof.
Corollary 3.11.
For the set $\text{DE}(X)$ of dependency equilibria of a game $X$, there are inclusions
$\overline{\Delta}\cap\overline{\mathcal{V}\setminus\mathcal{W}}^{\text{Zar}}=\overline{\Delta}\cap\overline{\mathcal{V}\setminus\mathcal{W}}\subseteq\text{DE}(G)\subseteq\overline{\Delta}\cap\mathcal{V},$
where $\mathcal{V}=\mathcal{V}_{X}$ is the Spohn variety of $X$. Both inclusions can be strict in general, see Section 4.
Remark 3.12.
Corollary 3.11 is still valid for generic games when one replaces complex tensors by real tensors in Definition 3.6(1). In other words, for generic games $X$, the following three statements are equivalent for a strategy $p$:
@@
For generic games, the Spohn variety $\mathcal{V}$ is irreducible and not contained in $\mathcal{W}$. Hence, (a) and (c) are equivalent. Moreover, the implication from (b) to (a) is trivial.
@@
In view of Corollary 3.11, it might be interesting to understand for which games the equality $\overline{\mathcal{V}\setminus\mathcal{W}}^{\text{Zar}}=\mathcal{V}$ holds. Indeed, in these cases, the set of DE coincides with $\overline{\Delta}\cap\mathcal{V}$. In general, $\overline{\mathcal{V}\setminus\mathcal{W}}^{\text{Zar}}=\mathcal{V}$ if and only if no irreducible component of $\mathcal{V}$ is contained in $\mathcal{W}$.
@@
The set $\mathcal{X}$ of all games $X\in\mathbb{R}^{n\times d_{1}\times\ldots\times d_{n}}$ such that $\overline{\mathcal{V}_{X}\setminus\mathcal{W}}^{\text{Zar}}=\mathcal{V}_{X}$ (which implies that the set of dependency equilibria of $X$ coincides with the non-negative real part of $\mathcal{V}_{X}$) is the complement of an affine variety in $\mathbb{R}^{n\times d_{1}\times\ldots\times d_{n}}$. In particular, it is a dense subset of $\mathbb{R}^{n\times d_{1}\times\ldots\times d_{n}}$.
@@
Problem 3.16.
For fixed $n$ and $d_{1},\ldots,d_{n}$, describe the Zariski open set $\mathcal{X}$ (see Prop. 3.15), that is, give the defining equations of its complement.
@@
Now, we want to give a first application of the lower and upper bounds for dependency equilibria from Corollary 3.11. It shows that, in reasonable situations, every Nash equilibrium is a dependency equilibria. By Example 4.7, although this is not always the case, by Theorem 3.18, every Nash equilibrium lies on the Spohn variety.
Note that the totally mixed Nash equilibria are exactly the positive real points that lie both on the Spohn variety and on the Segre variety by a result of Sturmfels and the first author [10, Theorem 6]. In particular, totally mixed Nash equilibria are always dependency equilibria.
@@
To recover the notation for general tensors on the Spohn variety, we set $p_{j_{1}\ldots j_{n}}=p_{j_{1}}^{(1)}\cdots p_{j_{n}}^{(n)}$. Note that $p_{+\ldots+k+\ldots+}=p_{k}^{(i)}$ for all $i$ and $k$. Let $i\in\{1,\ldots,n\}$ and $k,k^{\prime}\in\{1,\ldots,d_{i}\}$. If $p_{k}^{(i)}=0$ or $p_{k^{\prime}}^{(i)}=0$ then the corresponding defining equation for the Spohn variety is trivial. So, assume that both are non-zero. Then the corresponding defining equation of the Spohn variety boils down for rank-one tensors to (3), as noted in the proof of [10, Theorem 6]. ∎
@@
If the variety $\mathcal{W}$ contains no irreducible component of the Spohn variety $\mathcal{V}_{X}$ of a game $X$ then every Nash equilibrium of $X$ is a dependency equilibrium of $X$.
In particular, when the payoff tables of $X$ are generic, every Nash equilibrium of $X$ is a dependency equilibrium of $X$.
@@
Note that the pure strategy $p$ always lies on $\mathcal{V}$ by definition.
Proposition 3.21.
Suppose that the Spohn variety $\mathcal{V}$ of a game $X$ is smooth at the pure strategy $p$, that is, its Jacobian $\mathcal{J}(p)$ at $p$ (see Remark 3.20) has maximal possible rank. Then the following are equivalent
@@
    Every (euclidean) neighbourhood of $p$ contains a totally mixed dependency equilibrium of $X$.
@@
As $\mathcal{V}\cap U\cap\Delta\neq\emptyset$ for every neighbourhood $U$ of $p$, a whole irreducible component of $\mathcal{V}$ containing $p$ passes through $\Delta$ and is therefore not contained in $\mathcal{W}$. It follows by Corollary 3.11 that $p$ is a DE. ∎
@@
    Clearly, the method of the proof of Proposition 3.21 works for arbitrary $p\in\overline{\Delta}$ that is a smooth point of $\mathcal{V}$: Every euclidean neighbourhood of $p$ contains a DE of $X$ if and only if $\ker\mathcal{J}(p)+p$ contains a positive vector.
@@
    Even if this is not true, one might find out that $p$ is a DE of $X$ by checking that $\ker\mathcal{J}(p)+p$ is not contained in $\mathcal{W}$.
@@
    If $\ker\mathcal{J}(p)+p\subseteq\mathcal{W}$ it could still be possible to find out that $p$ is a DE of $X$ by exhibiting that an irreducible component of $\mathcal{V}$, containing $p$ as a smooth point, is not linear, for instance, by studying second order derivatives of $\mathcal{V}$ around $p$.
Even though the structure of $\mathcal{J}(p)$ is very simple for a pure strategy $p$, see Remark 3.20, which is helpful for concrete applications of Proposition 3.21, it is in general hard to derive a sensible general sufficient condition for the existence of DE around $p$ from it. However, it is possible for games of small size, such as $(2\times 2)$-games, see Example 4.8.
@@ 4. General results for $(2\times 2)$-games
Next, we exhibit when at least one of the four pure strategies is a smooth point on the Spohn variety of a game. Note that these points are always on the variety.
@@
    The converse of Theorem 3.10 is not true, that is, there might exist points of $\overline{\Delta}\cap\mathcal{V}$ that are no DE.
@@
Next, we give an example of a NE that is not a DE.
@@
In the following example, we work out and apply Proposition 3.21 in the case of $(2\times 2)$-games.
@@
Let us go back to the set-up of Proposition 4.1 and set $p=(1,0,0,0)$. In the situations (2) and (3), a vector $(w,x,y,z)$ in $\ker\mathcal{J}(p)$ must satisfy $z=0$ and hence, by Proposition 3.21, there is a neighbourhood of $p$ not containing a totally mixed DE of the game.
@@
So, these are exactly the cases in which every neighbourhood around $p$ contains totally mixed DE of the game by Proposition 3.21. In particular, if (4) is satisfied then $p=(1,0,0,0)$ is a pure DE of the game.
@@
We saw in Proposition 3.15 that the set of all games $X$ of a fixed size such that $\overline{\mathcal{V}_{X}\setminus\mathcal{W}}^{\text{Zar}}=\mathcal{V}$ (which, by Corollary 3.11, implies that $\mathcal{V}\cap\overline{\Delta}$ equals the set of DE of $X$) is the complement of an affine variety in the ambient affine space. In the following, we determine this set explicitly by its defining non-equations for $(2\times 2)$-games. As we will see, it is the complement of a finite union of linear subspaces of $\mathbb{C}^{8}$ and the restriction that it poses is also sensible in game-theoretic terms: It just asks for both players $i\in\{1,2\}$ that, if $i$ decides to choose a pure strategy, then $i$’s payoff is still dependent on the choice of pure strategy of the other player. This equals the usual notion of “generic $(2\times 2)$-game” within the game theory community and is made precise in (2) of the following theorem.
@@
In particular, if (2) holds then the set of dependency equilibria of $X$ equals the set of non-negative real points of $\mathcal{V}$ and every Nash equilibrium is a dependency equilibrium of $X$.
@@
The last sentence of the theorem follows from the equivalence above together with Corollary 3.11 and Corollary 3.19. So, it is left to prove the equivalence of (1) and (2).

### v3

@@ 3. Results for arbitrary games
Theorem 3.10 is still valid for generic games when one replaces complex tensors by real tensors in Definition 3.6(1). In other words, for generic games $X$, the following three statements are equivalent for a strategy $p$:
@@
For generic games, the Spohn variety $\mathcal{V}$ is irreducible and not contained in $\mathcal{W}$ by [10, Theorem 3.2] and Proposition 3.16. Hence, (a) and (c) are equivalent. Moreover, the implication from (b) to (a) is trivial.
@@
In view of Theorem 3.10, it might be interesting to understand for which games the equality $\overline{\mathcal{V}\setminus\mathcal{W}}^{\text{Zar}}=\mathcal{V}$ holds. Indeed, in these cases, the set of DE coincides with $\overline{\Delta}\cap\mathcal{V}$. In general, $\overline{\mathcal{V}\setminus\mathcal{W}}^{\text{Zar}}=\mathcal{V}$ if and only if no irreducible component of $\mathcal{V}$ is contained in $\mathcal{W}$.
@@
The set $\mathcal{X}$ of all games $X\in\mathbb{R}^{n\times d_{1}\times\ldots\times d_{n}}$ such that $\overline{\mathcal{V}_{X}\setminus\mathcal{W}}^{\text{Zar}}=\mathcal{V}_{X}$ (which implies that the set of DE of $X$ coincides with the non-negative real part of $\mathcal{V}_{X}$) is the complement of an affine variety in $\mathbb{R}^{n\times d_{1}\times\ldots\times d_{n}}$. In particular, it is a dense subset of $\mathbb{R}^{n\times d_{1}\times\ldots\times d_{n}}$.
@@
By Definition 2.2, every pure strategy lies on the Spohn variety. An immediate observation and an extension of [10, Theorem 3.2] is that every pure strategy is generically a DE.
Problem 3.17.
For fixed $n$ and $d_{1},\ldots,d_{n}$, describe the Zariski open set $\mathcal{X}$ (see Prop. 3.16), that is, give the defining equations of its complement.
@@
While Theorem 3.19 shows that every Nash equilibrium lies on the Spohn variety, Example 4.7 shows that not every Nash equilibrium is a DE. However, Corollary 3.20 gives a criterion for when a Nash equilibrium is a DE which is generically satisfied. Note that the totally mixed Nash equilibria are exactly the positive real points that lie both on the Spohn variety and on the Segre variety by a result of Sturmfels and the first author [10, Theorem 3.2]. In particular, totally mixed Nash equilibria are always DE.
@@
Figure 4. The real part of the Spohn variety consisting of four lines for the game with payoff tables $\begin{pmatrix}-1&-1\\ -2&3\end{pmatrix}\text{ and }\begin{pmatrix}-3&0\\ 0&0\end{pmatrix}$. The lower bound $\overline{\mathcal{V}\setminus\mathcal{W}}^{\text{Zar}}$ for DE equals the blue component. By Example 4.7, the blue point is a Nash equilibrium but not a dependency equilibrium. This is only possible because it lies on one of the green components of $\mathcal{V}$, see Corollary 3.20. These are exactly the components of $\mathcal{V}$ that are contained in $\mathcal{W}$.
@@
To recover the notation for general tensors on the Spohn variety, we set $p_{j_{1}\ldots j_{n}}=p_{j_{1}}^{(1)}\cdots p_{j_{n}}^{(n)}$. Note that $p_{+\ldots+k+\ldots+}=p_{k}^{(i)}$ for all $i$ and $k$. Let $i\in\{1,\ldots,n\}$ and $k,k^{\prime}\in\{1,\ldots,d_{i}\}$. If $p_{k}^{(i)}=0$ or $p_{k^{\prime}}^{(i)}=0$ then the corresponding defining equation for the Spohn variety is trivial. So, assume that both are non-zero. Then the corresponding defining equation of the Spohn variety boils down for rank-one tensors to (3), as noted in the proof of [10, Theorem 3.2]. ∎
@@
If the variety $\mathcal{W}$ contains no irreducible component of the Spohn variety $\mathcal{V}_{X}$ of a game $X$ then every Nash equilibrium of $X$ is a dependency equilibrium of $X$. In particular, when $X$ is generic, every Nash equilibrium of $X$ is a dependency equilibrium of $X$.
@@
Recall that the pure strategy $p$ always lies on $\mathcal{V}$ by definition.
Theorem 3.22.
Suppose that the Spohn variety $\mathcal{V}$ of a game $X$ is smooth at the pure strategy $p$, that is, its Jacobian $\mathcal{J}(p)$ at $p$ (see Remark 3.21) has maximal possible rank (the codimension of $\mathcal{V}$). Then the following are equivalent
@@
    Every (Euclidean) neighborhood of $p$ contains a totally mixed dependency equilibrium of $X$.
@@
As $\mathcal{V}\cap U\cap\Delta\neq\emptyset$ for every neighbourhood $U$ of $p$, a whole irreducible component of $\mathcal{V}$ containing $p$ passes through $\Delta$ and is therefore not contained in $\mathcal{W}$. It follows by Theorem 3.10 that $p$ is a DE. ∎
Based on Theorem 3.22, we will later give a very explicit form of this criterion for $(2\times 2)$-games involving only equalities and inequalities in the entries of the payoff tables, see Example 4.8. Figure 1 shows the Spohn variety of an example of a game where this criterion applies for the pure strategies $(1,0,0,0)$ and $(0,0,0,1)$, but it fails for the other two pure strategies, although these are also DE.
@@
    Clearly, the method of the proof of Theorem 3.22 works for arbitrary $p\in\overline{\Delta}$ that is a smooth point of $\mathcal{V}$: Every Euclidean neighborhood of $p$ contains a DE of $X$ if and only if $\ker\mathcal{J}(p)+p$ contains a positive vector.
@@
    What is more, one could still use the same ideas to show that $p$ is a DE of $X$ by checking that $\ker\mathcal{J}(p)+p$ is not contained in $\mathcal{W}$.
@@
    Even if $\ker\mathcal{J}(p)+p\subseteq\mathcal{W}$ it could still be possible to find out that $p$ is a DE of $X$ by exhibiting that an irreducible component of $\mathcal{V}$, containing $p$ as a smooth point, is not linear, for instance, by studying second order derivatives of $\mathcal{V}$ around $p$.
@@ 4. General results for $(2\times 2)$-games
Next, we show that at least one of the four pure strategies is always a smooth point on the Spohn variety of a game.
@@
$(2\times 2)$-games split up into distinct classes concerning the geometric structure of their sets of DE. An understanding of these classes is based on the considerations of Theorem 4.2 below and its proof which give, from this viewpoint, a full and sensible characterization of $(2\times 2)$-games. Cases (3)(c) and (3)(d) deserve further investigation concerning the number of irreducible components of the Spohn variety (which all are curves in these cases) and their individual degrees. This is studied in [8].
@@
    The converse of Lemma 3.9 is not true, that is, there might exist points of $\overline{\Delta}\cap\mathcal{V}$ that are not DE.
@@
Next, we give an example of a Nash equilibrium that is not a DE.
@@
In the following example, we work out and apply Theorem 3.22 in the case of $(2\times 2)$-games.
@@
Let us go back to the set-up of Proposition 4.1 and set $p=(1,0,0,0)$. In the situations (2) and (3), a vector $(w,x,y,z)$ in $\ker\mathcal{J}(p)$ must satisfy $z=0$ and hence, by Theorem 3.22, there is a neighbourhood of $p$ not containing a totally mixed DE of the game.
@@
So, these are exactly the cases in which every neighbourhood around $p$ contains totally mixed DE of the game by Theorem 3.22. In particular, if (4) is satisfied then $p=(1,0,0,0)$ is a pure DE of the game.
@@
We saw in Proposition 3.16 that the set of all games $X$ of a fixed size such that $\overline{\mathcal{V}_{X}\setminus\mathcal{W}}^{\text{Zar}}=\mathcal{V}$ (which, by Theorem 3.10, implies that $\mathcal{V}\cap\overline{\Delta}$ equals the set of DE of $X$) is the complement of an affine variety in the ambient affine space. In the following, we determine this set explicitly by its defining non-equations for $(2\times 2)$-games. As we will see, it is the complement of a finite union of linear subspaces of $\mathbb{C}^{8}$ and the restriction that it poses is also sensible in game-theoretic terms: It just asks for both players $i\in\{1,2\}$ that, if $i$ decides to choose a pure strategy, then $i$’s payoff is still dependent on the choice of pure strategy of the other player. This equals the usual notion of “generic $(2\times 2)$-game” within the game theory community and is made precise in (2) of the following theorem.
@@
In particular, if (2) holds then the set of DE of $X$ equals the set of non-negative real points of $\mathcal{V}$ and every Nash equilibrium is a dependency equilibrium of $X$.
@@
The last sentence of the theorem follows from the equivalence above together with Theorem 3.10 and Corollary 3.20. So, it is left to prove the equivalence of (1) and (2).